OpenStudy (anonymous):

help me with this integration : 1/2 \[\int\limits_{?}^{?} (4x + 6)/\sqrt{2x^2 +6x -1}\] dx

OpenStudy (mimi_x3):

What's the question mark>

OpenStudy (anonymous):

there no limit..leave the answer in terms of x

OpenStudy (mimi_x3):

Let u = 2x^2 + 6x - 1 du/ 4x + 6

OpenStudy (binary3i):

\[2\sqrt{2x^2+6x-1}\]

OpenStudy (mimi_x3):

\[\int\limits\frac{4x+6}{\sqrt{2x^{2}+6x-1}} \] Let u = 2x^2 +6x - 1 \[\int\limits\frac{4x+6}{\sqrt{u}} *\frac{du}{4x+6} \] \[\int\limits\frac{1}{\sqrt{u}}du = \int\limits\frac{u^{1//2}}{1/2} \] \[\frac{(2x^{2}+6x-1)^{1/2}}{1/2} + C\]

OpenStudy (mathmate):

Note that the original question has a factor (1/2) written above the integral. If that's the case, both results have to be divided by 2.

OpenStudy (mimi_x3):

I dont think that the (1/2) is related to the integral :/

OpenStudy (mathmate):

Perhaps, but it can also happen when the 1/2 was typed in before entering the "equation" button, which makes the integral go to a separate line. In any case, the results correspond clearly to the integral without the (1/2).

OpenStudy (mimi_x3):

Ohh woops, my bad i misunderstood the question, sorry. I thought the (1/2) was nothing xD